OpenStudy (anonymous):

what is the simplest form of the expression 5(x+2)_ 7(x+2)

OpenStudy (anonymous):

5(x+2)= 5*x+2*5=5x+10 7(x+2)=7*x+7*2=7x+14

OpenStudy (anonymous):

u need more explanation ?

OpenStudy (anonymous):

not really thx so much

OpenStudy (anonymous):

your welcome :)